LCG Service Coordination Meeting

→ Europe/Zurich
513/1-024 (CERN)

513/1-024

CERN

50
Show room on map
Jamie Shiers
Description
LCGSCM

Internal meeting for host laboratory LCG service providers focussing on issues on a timescale of ~1 week to ~1 month.

Mailing list: lcg-service-coordination-meeting@cern.ch

Present: J.Shiers (chairman), H.Renshall (secretary), J.Casey, G.Mccance, O.Barring, T.Kleinwort, A.Retico, S.K, U.Schwickerath, J.v.Eldik, P.Mendez, V.Boppanna, V.Singhal, M.Anjo, M.Girone, S.Campana, R.Santinelli, A.Sciaba, M.Dimou, R.Mollon, M.Littmaath Outstanding Issues and Actions: Reinstallation: TK- ongoing. Apel (grid accounting) problems at the moment. Hardware for SLC4 tests: D.Smith has requested 4 boxes (ex-batch nodes- pre-agreed). CMS Servers: HR- we should revive IT-CMS task force. I will contact CMS. Other changes: JvE- have to upgrade firmware of 118 Transtec diskservers in February. Requires node downtime. MG- We need to schedule expansion of RAC clusters. Should be less than 4 hours. Highlevel milestones: As attached. JS- agreed at yesterdays MB meeting. LCG Service Review: Certification: As attached. Monitoring: JC- have had 35 replies to survey of site monitoring tools. WMS and DM: As attached. No news on the grid UI for SLC4 but the lxplus alias change to the SLC4 cluster will not be delayed. Databases: As attached. Auth and Auth: As attached. LHCB: RS- see attached and FZK problem was fixed today (an NFS mount of lhcb software) ALICE: PM- see attached and wanting to develop experiment SAM tests. ATLAS: SC- Xmas MC production was best ever. Main problem is voms production role cannot write in user DPM directories. Have script from J-P.Baud to change ACLs and would like to test it on PPS but we will ask for long term solution of allowing secondary groups in DPM and LFC. We are also interested in SAM tests for the site vo boxes. CMS: AS- we are testing SLC4 worker nodes. We had a problem with users disappearing from voms because their institute expired in the CERN HR database. AOB: TK- L.Field wants to do some bdii testing. T.Bell has discovered some older lfc and bdii nodes were installed as jbod system disks rather than mirrored. JS- during the WLCG workshop next week please participate in the panel sessions to try and keep them lively and relevant.
There are minutes attached to this event. Show them.
    • 10:00 → 10:10
      Outstanding Issues & Actions 10m
      • Reinstallation of service nodes - target end 2006
      • H/W availability for testing 64bit / SLC4
      • CMS servers / services: status
        See CMS task force of Sep 20
      • Other changes in the pipeline
      • WLCG High Level Milestones for 2007
        Slides
      • Input to CERN report to WLCG Collaboration workshop next week 15m
      • <a href="https://twiki.cern.ch/twiki/bin/view/LCG/LcgScmStatusActions">Action Items - Summary</a>
        Speaker: Filled in at the meeting
    • 10:10 → 10:40
      LCG Service Review 30m
      • <a href="https://twiki.cern.ch/twiki/bin/view/LCG/LcgScmStatusDeploy">Certification / Pre-production</a>
        Speakers: Nick Thackray, Oliver Keeble
        PPS status
      • <a href="https://twiki.cern.ch/twiki/bin/view/LCG/LcgScmStatusMLR">Monitoring, Logging & Reporting</a>
        Speakers: Ian Neilson, James Casey
      • Core Grid Services
        Speakers: Gavin McCance, Jan van Eldik, Thorsten Kleinwort, Ulrich Schwickerath
      • Fabric & Infrastructure Services
        Speakers: Maria Dimou, Maria Girone, Remi Mollon
      • Other Service Issues
        Speaker: GMOD, SMOD
      • <a href="https://twiki.cern.ch/twiki/bin/view/LCG/LcgScmStatusExp">Experiment Issues</a>
        Speaker: EIS Team
    • 10:40 → 10:45
      Any Other Business 5m